# Active Inference for Self-Organizing Multi-LLM Systems: A Bayesian Thermodynamic Approach to Adaptation

### v1

@@ 4.4 Action Selection
The agent employs a sophisticated policy selection mechanism with a policy horizon of 2 steps and an inference horizon of 1 step. Both state-information gain and parameter-information gain are enabled, and the model uses deterministic action selection. Valid policies are constructed to allow three types of actions: no action [0 0 0], prompt changes only [¿0 0 0], and search actions only [0 ¿0 0].
@@ 5.1 Learning Environment Dynamics
Through active exploration and learning, the agent successfully developed an accurate model of the environment dynamics, particularly the relationships between states and observations. Figure 2 shows the final learned observation mappings after multiple interactions with the environment. Compared to the initial uniform distributions (Figure 1), these matrices show clear structure, indicating the agent has learned meaningful relationships between states and observations. The prompt quality matrices ($A_{[0,1,2]}$) developed distinct patterns showing which prompt combinations lead to higher quality outputs. The search quality matrices ($A_{[3,4,5]}$) reveal learned associations between search actions and information quality, while the information state matrix ($A_{[6]}$) captures the reliable mapping between hidden and observed information states.
@@ 5.2 Strategic Action Selection
The agent’s action selection strategy evolved over time, demonstrating increasingly sophisticated decision-making. Figure 3 shows the expected free energy of different policies at four time points during the agent’s operation. This progression reveals how the agent learned to value different action combinations based on their information-gathering and goal-achieving potential.
@@
5.3 Action Selection Patterns
The agent’s actual behavior patterns provide strong evidence for intelligent exploration and exploitation. Figure 4 shows the frequency of different action combinations throughout the experimental runs. The distribution of actions reveals a strategic approach to environment exploration.
Figure 4: Heatmap showing the frequency of action selection across prompt and search dimensions. Darker colors indicate more frequently selected actions. The pattern shows strategic exploration focusing on information-gathering search actions early in the process, followed by targeted prompt testing informed by acquired knowledge.
@@ 5.4 Information-Driven Exploration
A key finding is the agent’s strategic use of search actions to minimize entropy and guide prompt exploration. The action selection patterns show that the agent learned to prioritize information-gathering search actions when uncertainty was high, using the resulting knowledge to inform subsequent prompt testing. This behavior emerged naturally from the free energy minimization framework, with the agent recognizing that search actions could reduce uncertainty about the environment’s structure.
Early in the learning process, the agent showed a higher proportion of search actions, as evidenced by the concentration of activity in the search action dimension of Figure 4. As the agent gathered more information and reduced uncertainty about the environment, it transitioned to more focused prompt testing, demonstrating an effective balance between exploration and exploitation.
@@
The progression of policy EFE values (Figure 3) shows how the agent’s evaluation of different action combinations became more refined over time. Initially, the differences in EFE values were relatively small, reflecting high uncertainty about action outcomes. As learning progressed, the EFE landscape developed clear structure, indicating the agent had learned which action combinations were most effective for achieving its goals.
@@
The final learned observation matrices (Figure 2) provide clear evidence that this exploration strategy was successful in discovering the underlying structure of the environment. The emergence of distinct patterns in these matrices, particularly in the prompt-quality relationships, demonstrates that the agent effectively learned which prompt combinations lead to better outcomes.

### v2

@@ 4.3.2 Transition Learning ($pB$)
4.4 Observation Generation
The agent receives structured observations about search quality and prompt effectiveness through evaluations performed by GPT-4o-mini, which provides standardized JSON output. For example, for search-related observations, the model evaluates search results using a structured output format:
Return only a JSON object with these three metrics, scored 0.0-1.0:
    "info_relevance": [Score from 0.0 to 1.0 counting up by 0.1],
    "info_usefulness": [Score from 0.0 to 1.0 counting up by 0.1],
    "source_quality": [Score from 0.0 to 1.0 counting up by 0.1]
These standardized scores are then scaled to the 11-point observation space used by the active inference agent’s observation model. The use of structured JSON output ensures consistent, programmatic evaluation of search results, providing reliable feedback for the agent’s learning process. Similar structured evaluations are performed for prompt quality metrics (accuracy, relevance, and comprehensiveness), with the language model providing standardized scores that are then mapped to the agent’s observation space.
This approach leverages GPT-4o-mini’s capability for structured output generation to create a reliable evaluation pipeline, ensuring that the active inference agent receives consistent, well-formatted observations for updating its belief states and learning the effectiveness of different actions.
@@ 4.5 Action Selection
The agent employs a sophisticated policy selection mechanism with a policy horizon of 2 steps and an inference horizon of 1 step. Both state-information gain and parameter-information gain are enabled, and the model uses deterministic action selection. Valid policies consist of three types of actions: no action $[0,0,0]$, prompt-only actions $[p,0,0]$ where $p\in\{1,2,\ldots,33\}$, and search-only actions $[0,s,0]$ where $s\in\{1,2,\ldots,11\}$. There can be no such action where $[p,s,0]$ since this would imply that the prompt and search actions are being performed simultaneously and the agent is not able to do that.
@@ 5.1 Learning Environment Dynamics
Through active exploration and learning, the agent successfully developed an accurate model of the environment, particularly the relationships between states and observations. Figure 2 shows the final learned observation mappings after multiple interactions with the environment. Compared to the initial uniform distributions (Figure 1), these matrices show clear structure, indicating the agent has learned meaningful relationships between states and observations. The prompt quality matrices ($A_{[0,1,2]}$) developed distinct patterns showing which prompt combinations lead to higher quality outputs. The search quality matrices ($A_{[3,4,5]}$) reveal learned associations between search actions and information quality, while the information state matrix ($A_{[6]}$) captures the reliable mapping between hidden and observed information states.
@@ 5.2 Strategic Action Selection
The agent’s action selection strategy evolved over time as its Expected Free Energy over policies changed. Figure 3 shows the expected free energy of different policies at four time points during the agent’s operation. This progression reveals how the agent learned to value different action combinations based on their information-gathering and goal-achieving potential.
@@ 5.3 Information-Driven Exploration
The agent’s action selection patterns provide strong evidence for sophisticated exploration and exploitation behavior emerging from the free energy minimization framework. Figures 4 and 5 illustrate complementary views of this behavioral progression.
Figure 4: Heatmap showing the frequency of action selection across prompt and search dimensions. Lighter colors indicate more frequently selected actions. The pattern shows the overall distribution of action selections, with certain prompt-search combinations being consistently preferred over others based on their effectiveness.
Figure 5: Time series of action selection throughout the experiment. Blue dots represent prompt actions (labeled with prompt IDs), while red dots represent search actions (labeled with search IDs). The progression shows a clear transition from search-dominated early phases to prompt-dominated later phases, demonstrating the agent’s evolving strategy from exploration to exploitation.
The temporal progression of the agent’s behavior, clearly visible in Figure 5, reveals a strategic shift in action selection. During the initial phase (approximately the first 40 timesteps), the agent heavily favors search actions, indicated by the higher density of red dots. This initial search-focused behavior naturally emerges from the free energy minimization framework, as the agent prioritizes reducing uncertainty about its environment through observation of search quality metrics and information state levels.
As the experiment progresses, Figure 5 shows a marked transition to prompt-dominated behavior, evidenced by the increasing density of blue dots in later timesteps. This shift occurs as the agent receives observations indicating higher information states and favorable search quality metrics. Rather than directly encoding search results, the agent’s behavior is guided by these external indicators of knowledge state and search effectiveness, leading to more focused prompt testing.
The aggregate view provided by Figure 4 complements this temporal analysis by showing the overall distribution of action selections. The concentration patterns in the heatmap reveal which prompt-search combinations proved most effective, where prompt and search actions that produced more preferred observations were sampled more.
@@
The progression of policy Expected Free Energy (EFE) values (Figure 3) provides further insight into how the agent’s evaluation of different action combinations became more refined over time. Initially, the differences in EFE values were relatively small, reflecting the agent’s inability to differentiate between the available policies. As learning progressed, the EFE landscape developed clear structure, indicating the agent had learned which action combinations were most effective for achieving its goals.
@@
The final learned observation matrices (Figure 2) provide evidence that this exploration strategy was successful in discovering the underlying structure of the environment. The emergence of distinct patterns in these matrices, particularly in the prompt-quality relationships, demonstrates that the agent effectively learned which prompt combinations lead to better outcomes, guided by the externally provided information state observations and quality metrics rather than direct knowledge incorporation from searches.
This sophisticated exploration-exploitation pattern emerged naturally from the free energy minimization framework, with the agent learning to balance information-gathering (through search actions and their associated quality metrics) with exploitation of known effective prompts (guided by prompt quality observations and information state feedback). The clear temporal progression from exploration to exploitation, visible in both the timeline and aggregate statistics, demonstrates the effectiveness of active inference in managing this fundamental trade-off.
6 Future Work
This paper has demonstrated a novel approach to combining active inference with traditional AI agents, showing how active inference can guide the exploration and optimization of agent parameters. However, this implementation represents only an initial step toward fully adaptive AI systems.
A key limitation of the current model is its reliance on a fixed state space with predefined prompts. Future work should focus on developing active inference models capable of dynamically expanding their state space during operation. This would allow the agent to explore a vastly larger set of possible prompts and configurations, moving beyond the constraints of preset options to discover novel, potentially more effective combinations.
Another crucial direction is the development of hierarchical generative models. Rather than relying on abstract information states, such models would enable the agent to directly encode and reason about environmental information. This would allow for more sophisticated understanding of the agent’s performance and context, leading to more informed adaptation decisions.
The framework could also be extended to optimize broader aspects of agent intelligence, including agent architecture, execution order of sub-agents, and tool utilization. These elements represent critical parameters that require intelligence to optimize effectively. By incorporating these factors into the active inference model, we could enable more comprehensive agent improvement.
These developments represent intermediate steps toward an ambitious long-term goal: a universal active inference model that can be applied to any agent system without predefined parameters. Such a model would function analogously to human researchers, capable of conducting research, engaging in trial and error, performing comprehensive parameter testing, and maintaining an intelligent exploration-exploitation loop. This would enable automated, principled improvement of any given agent system, representing a significant advance in artificial intelligence adaptation and optimization.
Data Availability
The full code implementation of the active inference agent, environment, and evaluation system described in this paper is available in our public GitHub repository at
https://github.com/RPD123-byte/Active-Inference-for-Self-Organizing-Multi-LLM-Systems-A-Bayesian-Thermodynamic-Approach-to-Adaptat.
This includes all components necessary to reproduce our experimental results, including the agent controller, environment simulation, and visualization tools.